# Refining Pre-Trained Motion Models

### 01 Jan 2024 v1

@@ IV-E Ablation Studies
TABLE V: Ablation study in MPI-Sintel-Clean. Results are evaluated on $8$ videos from the Sintel-Clean train split. Both Cycle Consistency and Color Constancy self-supervised metrics lead to improvements from the pretrained RAFT. Combining both metrics works the best.
@@ V Discussion and conclusion
In this paper, we present a procedure to improve state-of-the-art supervised motion models with only self-supervised fine-tuning. Our procedure is a two-stage method where we first use a pretrained motion model to sample and filter cycle-consistent tracks in the test domain, and then we fine-tune on these tracks, with added augmentations. We compare our method with state-of-the-art supervised models, self-supervised training methods, and self-supervised finetuning techniques. We find our method improves the pretrained model performance, on three different datasets, with two different tracking models, whereas existing self-supervised techniques worsen performance. We hope our method will inspire future work on refining pre-trained motion models.

### 17 Feb 2024 v2

@@ IV-E Ablation Studies
TABLE V: Ablation study in MPI-Sintel-Clean. Results are evaluated on $8$ videos from the Sintel-Clean train split. Both Cycle Consistency and Color Constancy self-supervised metrics lead to improvements from the pre-trained RAFT. Combining both metrics works the best.
@@ V Discussion and conclusion
In this paper, we present a procedure to improve state-of-the-art supervised motion models with only self-supervised fine-tuning. Our procedure is a two-stage method where we first use a pre-trained motion model to sample and filter cycle-consistent tracks in the test domain, and then we fine-tune on these tracks, with added augmentations. We compare our method with state-of-the-art supervised models, self-supervised training methods, and self-supervised finetuning techniques. We find our method improves the pre-trained model performance, on three different datasets, with two different tracking models, whereas existing self-supervised techniques worsen performance. We hope our method will inspire future work on refining pre-trained motion models.
VI Acknowledgments.
This work was supported by the Toyota Research Institute under the University 2.0 program, ARL grant W911NF-21-2-0104, and a Vannevar Bush Faculty Fellowship. The authors thank Leonid Keselman for suggesting the title.